OpenStudy (anonymous):

5vC3v I really don't have any idea how to solve this equation. Look at the first post to see what is looks like on my paper.

OpenStudy (anonymous):

\[_{5}C _{3}\]

OpenStudy (accessdenied):

\[ \large _{n}C_{k} = \frac{n!}{k! (n - k)!} \]

OpenStudy (accessdenied):

Which we use n = 5 and k = 3 in that expression to evaluate it

OpenStudy (anonymous):

1.2?

OpenStudy (accessdenied):

No, these should come out to be integers. If we write out the factorial: \[ \frac{5\times4\times3\times2\times1}{(3\times2\times 1)(2\times1)}\] There, we can just cancel out the things on the numerator and denominator. 3/3, 2/2, and 4/2 = 2/1

OpenStudy (accessdenied):

\[ \frac{5\times\cancel{4}2\times \cancel{3}\times\cancel{2}\times1}{(\cancel{3}\times\cancel{2}\times 1)(\cancel{2}\times1)} \]

OpenStudy (accessdenied):

so that just leaves 5*2 = 10

OpenStudy (anonymous):

yeah, 10 thanks :D Best answer, Closed. thanks again @AccessDenied !

OpenStudy (accessdenied):

You're welcome. :)